# Black Hole Entropy, Quantum Corrections and EFT Transitions

### v1

@@ 2.2 An exact entropy formula for BPS black holes
The physical significance of $\mathscr{Z}$ is that it controls the warp factor of the metric in the BPS black hole background, whose near-horizon line element reads (using isotropic coordinates) Ferrara:1997yr ; LopesCardoso:1998tkj
$ds^{2}=-e^{2U(r)}dt^{2}+e^{-2U(r)}\left(dr^{2}+r^{2}d\Omega_{2}^{2}\right)\,,\qquad\text{with}\ \ e^{-2U(r)}=1+\frac{|\mathscr{Z}|^{2}\kappa_{4}^{2}}{8\pi r^{2}}\,,$ (2.19)
thus also incorporating the effect of the higher-derivative chiral terms captured by eq. (2.5). The attractor equations then determine the values for the moduli fields $Y^{A}$ when evaluated at the horizon to be fixed by Behrndt:1998eq
@@
We conclude this section by giving some details on the quantum entropy formula presented above. The non-interested reader can safely skip this discussion. First of all, let us note that (2.21) has been computed using Wald’s prescription Wald:1993nt ; Iyer:1994ys within the restricted framework of conformal off-shell $\mathcal{N}=2$ supergravity coupled to $n_{V}+1$ vector multiplets Ferrara:1977ij ; deWit:1979dzm ; deWit:1980lyi ; deWit:1984rvr ; deWit:1984wbb , which reduces to the more familiar 4d $\mathcal{N}=2$ (Poincaré) supergravity only after partial gauge fixing. [^5] The relation between conformal and Poincaré (extended) supegravity requires the introduction of an additional vector multiplet that can be used to gauge-fix dilatation invariance deWit:1980lyi . See also VanProeyen:1983wk ; deWit:1984hw for early reviews on the topic. This formalism can be used, in turn, to derive the attractor equations (2.20) as well as the near-horizon metric (2.19). This means, consequently, that (2.21) provides the macroscopic entropy associated to BPS black hole solutions in Calabi–Yau compactifications of Type IIA string theory, when restricting ourselves to the gravity and vector multiplet sectors. [^6] As is well known, the two-derivative theory can be truncated consistently. The quantum corrections to the hypermultiplet sector, on the other hand, are not fully known Antoniadis:1993ze ; Robles-Llana:2006vct and thus we cannot verify whether higher-derivative terms involving those will obstruct the truncation. Notice, however, that upon doing so we might be missing some contributions due to non-chiral higher-derivative operators (i.e., those intrinsically defined as integrals over full superspace) in the vector-multiplet sector, as well as analogous hypermultiplet-dependent terms in the 4d effective action. A large class of the former type of couplings were already shown to give a vanishing contribution to the black hole entropy deWit:2010za ; Murthy:2013xpa , hence suggesting that this could always be the case. As for the latter, in LopesCardoso:2000qm it was explicitly checked that adding neutral hypermultiplets in the form of gauge-fixed, superconformal multiplets deWit:1999fp does not affect neither the attractor mechanism, nor the BPS near-horizon geometry. The analysis therein was carried out by considering perturbative $\mathcal{R}^{2}$-corrections. However, there is no guarantee that this will still work at all orders in perturbation theory. In fact, the authors of Ooguri:2004zv argue —also providing some amount of evidence— that the exact black hole entropy should depend on the background hypermultiplet vevs. Crucially, though, the generalized prepotential (2.18) controlling the quantum entropy formula above is sensitive to the number of hypermultiplets but not to their vevs (see discussion after eq. (2.29) in the next section). Thus, from the macroscopic perspective it is not clear whether we could be missing some additional operators contributing to the black hole entropy, namely if (2.21) would be the end result of applying Wald’s procedure in the full Type IIA string theory. In Ooguri:2004zv a detailed analysis of the origin of (2.21) was performed and they suggested that it is computing instead a protected supersymmetric index. This idea has been supported by explicitly matching the black hole free energy [^7] This is the Legendre dual of the entropy and it is the leading contribution to the gravitational path integral. with a supersymmetric index defined within the CFT living on the branes sourcing the BPS black hole background. In particular, the alternating signs of the terms which add up to give the index should account for the cancellation of the dependence on hypermultiplet vevs of the BPS states degeneracy. In what follows, we will not be concerned about whether (2.21) is truly computing an entropy or a protected supersymmetric index in Type IIA, and we will just focus on its properties along certain decompactification limits. With this subtlety in mind, we will refer to (2.21) simply as the BPS black hole entropy.
@@
It is also worth mentioning that in Ooguri:2004zv they revise the relation between macroscopic entropy and microstate counting performed in Strominger:1996sh . We present the subtlety following the modern review of Zaffaroni:2019dhb (see also references therein). What one should truly compute in order to compare the macroscopic and microscopic dual descriptions of the system is the partition function $\mathcal{Z}$. Such quantity can be formally defined via some path integral or as a microscopic generating function, respectively. In the latter case, it is not defined with a micro-canonical ensemble (i.e., with both electric $q$ and magnetic $p$ charges fixed), but rather with a mixed ensemble (fixed magnetic charges $p$ and electric potentials $\phi$). Thus, for a supersymmetric system, such partition function would have the form [^8] In the general case, there would also be some dependence on the angular momentum, which we omit here.
@@
with $F$ being some $\mathbb{Z}_{2}$-graded (i.e., fermionic) operator. The advantage of using the index is that it can also be evaluated from the macroscopic perspective. It is indeed an euclidean path integral with proper boundary conditions (which can be explicitly determined in concrete examples). In general, a supersymmetric index does not coincide with the partition function. However, in particular setups one can actually prove that the index provides a good approximation to the partition function. In essence, what one has to ensure is that there are no large cancellations between the different supersymmetric states over which we are tracing. [^10] In some examples this is automatically realized thanks to the symmetries exhibited by the configuration. In other case, the cancellation is avoided if the chemical potentials involved have complex phases. In general, though, there is not a unique, unambiguous prescription to find an appropriate $(-1)^{F}$ operator. Then, what Ooguri:2004zv suggests is that, despite (2.21) not being constructed as an index, what it is truly computing in the context of Type IIA is the Legendre dual of $\mathcal{Z}_{\text{index}}$, and is therefore a protected quantity. Moreover, in the large charge expansion, we would also have $\mathcal{Z}_{\text{index}}\sim\mathcal{Z}$, so that it really computes the same quantity as $\mathcal{S}_{\text{micro}}$ and $\mathcal{S}_{\text{BH}}$, cf. eqs. (2.23) and (2.24).
@@ 2.3.1 Leading-order corrections to the generalized prepotential
$\mathcal{F}(X^{A})=$ $-\frac{1}{6}\mathcal{K}_{abc}\frac{X^{a}X^{b}X^{c}}{X^{0}}+K_{ab}^{(1)}X^{a}X^{b}+K_{a}^{(2)}X^{0}X^{a}+K^{(3)}$ (2.26)
@@
$\mathcal{K}_{abc}=\omega_{a}\cdot\omega_{b}\cdot\omega_{c}\,,\qquad K_{a}^{(2)}=\frac{1}{24}c_{2}(X_{3})\cdot\omega_{a}\,,\qquad K^{(3)}=\frac{i\zeta(3)}{2(2\pi)^{3}}\chi_{E}(X_{3})\,,$ (2.27)
whereas $K_{ab}^{(1)}$ can be fixed instead by requiring good symplectic transformation properties of the underlying period vector deWit:1992wf ; Harvey:1995fq . Similarly, the quantity $\chi_{E}(X_{3})=2\left(h^{1,1}(X_{3})-h^{2,1}(X_{3})\right)$ denotes the Euler characteristic of the threefold. Lastly, the coefficients $n_{\boldsymbol{k}}^{(0)}$ are known as genus-zero Gopakumar–Vafa invariants and count, for each positive homology representative $\boldsymbol{k}=k_{a}\gamma^{a}\in H_{2}^{+}(X_{3},\mathbb{Z})$, the indexed degeneracy of supersymmetric D2-brane states wrapped on 2-cycles within the corresponding holomorphic 2-cycle class Gopakumar:1998ii ; Gopakumar:1998jq .
On the other hand, the genus-1 topological string amplitude can be expanded around the large radius point as Bershadsky:1993ta ; Bershadsky:1993cx ; Katz:1999xq
$\mathcal{F}_{1}(X^{A})=\frac{1}{24}\int_{X_{3}}J_{\mathbb{C}}\wedge c_{2}(TX_{3})\,+\,\mathcal{O}(\log z^{a})=\frac{1}{24}c_{2,a}\,z^{a}\,+\,\mathcal{O}\left(e^{2\pi iz^{a}}\right)\,,$ (2.28)
where $J_{\mathbb{C}}=z^{a}\,\omega_{a}=(b^{a}+it^{a})\,\omega_{a}$ is the complexified Kähler 2-form of the Calabi–Yau threefold, whilst $c_{2}(TX_{3})$ denotes the second Chern class of its tangent bundle. This contribution can be easily understood as coming from the dimensional reduction of the analogous $\text{tr}\,\mathcal{R}\wedge\mathcal{R}$ operator in 5d $\mathcal{N}=1$ supergravity Grimm:2017okk .
@@
where $n\in\mathbb{Z}$ is the 0-brane charge. Therefore, upon substituting this into (2.10) and performing the integral —taking account that each D0-brane yields $-\frac{\chi(X_{3})}{2}$ times the contribution of a single hypermultiplet Gopakumar:1998jq — as well as the infinite sum, one finds (in units of $m_{\rm D0}/2\pi$)
@@ 2.3.2 Black hole solutions in the large volume patch
[…]
@@
and we isolate in (2.39) the terms depending only on $G(Y^{0},\Upsilon)$ together with its derivatives, we obtain the simpler formula
@@ 3.1.1 The two-derivative analysis
with $C=e^{K/2}\bar{Z}$ some compensator field that ensures the symplectic invariance of any solution to the attractor equations above (cf. discussion around (2.18)). Henceforth, we will concentrate on black holes characterized by having $p^{0}=0$, i.e., no D6-brane charge, as seen from the Type IIA perspective. Consequently, we deduce from (3.3) that the rescaled quantity $CX^{0}$ is purely real (and, e.g., non-negative), hence allowing us to completely solve the algebraic system as follows
@@
where $D^{ab}$ is the inverse matrix of $D_{ab}\equiv D_{abc}p^{c}$ —which is assumed to exist, [^11] This ensures that there exists a unique solution to the algebraic system (3.3), given precisely by (3.4). Notice that $D_{ab}$ could be singular in special circumstances, such that $D^{ab}$ might not be well-defined Marchesano:2023thx ; Marchesano:2024tod ; Castellano:2024gwi ; CMP . and $\hat{q}_{0}\equiv q_{0}+\frac{1}{12}D^{ab}q_{a}q_{b}$. Note that this latter shift may be interpreted as an induced D0-brane charge in the worldvolume of the D2- and D4-branes comprising the 4d black hole of the form $\delta q_{0}=q_{a}\,\text{Re}\,z^{a}-3p^{a}D_{abc}\,\text{Re}\,z^{b}\,\text{Re}\,z^{c}$.
@@
$(\hat{q}^{0})^{2},\,\left(p^{a}\right)^{2}\gg\left|\frac{D_{abc}p^{a}p^{b}p^{c}}{\hat{q}_{0}}\right|\,.$ (3.9)
@@ 3.2.1 Including higher-derivative corrections
[…]
@@
Finally, let us try to understand the extra constraints imposed by the hierarchy (3.12). The first one is required so as to ensure that the corrections to the attractor solution and black hole entropy due to the genus-1 term in (2.31) are in fact subleading with respect to the two-derivative results. The second condition, on the other hand, is more interesting, since its net effect is to suppress the higher-genus terms as well. [^12] Actually, it suppresses the full tower of quantum corrections $\delta\mathcal{F}_{g}$ associated to D0-brane states, as captured by $G(Y^{0},\Upsilon)$, which also includes a genus-0 contribution. Furthermore, it can be translated into an equivalent mathematical statement on the value for $y^{0}$ (equivalently $x^{0}$ in (3.4)), whose asymptotics have been ignored so far. To see this, we should first compute the imaginary part of $G_{0}(Y^{0},\Upsilon)$, which from the local 4d EFT perspective is given by the following asymptotic series
@@
$(\hat{q}^{0})^{2},\,\left(p^{a}\right)^{2}\gg\left|\frac{D_{abc}p^{a}p^{b}p^{c}}{\hat{q}_{0}}\right|\gg 1\,,\qquad\text{with}\quad|\hat{q}^{0}|\,,\,p^{a}\gg 1\,.$ (3.18)

### v2

@@ 2.2 An exact entropy formula for BPS black holes
Physically, the quantity $\mathscr{Z}$ controls the warp factor of the metric in the BPS black hole background Ferrara:1997yr ; LopesCardoso:1998tkj , whose near-horizon line element reads (using isotropic coordinates)
$ds^{2}=-e^{2U(r)}dt^{2}+e^{-2U(r)}\left(dr^{2}+r^{2}d\Omega_{2}^{2}\right)\,,\qquad\text{with}\ \ e^{-2U(r)}=\frac{|\mathscr{Z}|^{2}\kappa_{4}^{2}}{8\pi r^{2}}\,$, (2.19)
thus also incorporating the effect of the higher-derivative chiral terms captured by eq. (2.5). The attractor equations then determine the values for the moduli fields $Y^{A}$ when evaluated at the horizon to be fixed by Behrndt:1998eq ; LopesCardoso:2000qm
@@
We conclude this section by giving some details on the quantum entropy formula presented above. The non-interested reader can safely skip this discussion. First of all, let us note that (2.21) has been computed using Wald’s prescription Wald:1993nt ; Iyer:1994ys within the restricted framework of conformal off-shell $\mathcal{N}=2$ supergravity coupled to $n_{V}+1$ vector multiplets Ferrara:1977ij ; deWit:1979dzm ; deWit:1980lyi ; deWit:1984rvr ; deWit:1984wbb , which reduces to the more familiar 4d $\mathcal{N}=2$ (Poincaré) supergravity only after partial gauge fixing. [^5] The relation between conformal and Poincaré (extended) supegravity requires the introduction of an additional vector multiplet that can be used to gauge-fix dilatation invariance deWit:1980lyi . See also VanProeyen:1983wk ; deWit:1984hw for early reviews on the topic. This formalism can be used, in turn, to derive the attractor equations (2.20) as well as the near-horizon metric (2.19). This means, consequently, that (2.21) provides the macroscopic entropy associated to BPS black hole solutions in Calabi–Yau compactifications of Type IIA string theory, when restricting ourselves to the gravity and vector multiplet sectors. [^6] As is well known, the two-derivative theory can be truncated consistently. The quantum corrections to the hypermultiplet sector, on the other hand, are not fully known Antoniadis:1993ze ; Robles-Llana:2006vct and thus we cannot verify whether higher-derivative terms involving those will obstruct the truncation. Notice, however, that upon doing so we might be missing some contributions due to non-chiral higher-derivative operators (i.e., those intrinsically defined as integrals over full superspace) in the vector-multiplet sector, as well as analogous hypermultiplet-dependent terms in the 4d effective action. A large class of the former type of couplings were already shown to give a vanishing contribution to the black hole entropy deWit:2010za ; Murthy:2013xpa , hence suggesting that this could always be the case. As for the latter, in LopesCardoso:2000qm it was explicitly checked that adding neutral hypermultiplets in the form of gauge-fixed, superconformal multiplets deWit:1999fp does not affect neither the attractor mechanism, nor the BPS near-horizon geometry. The analysis therein was carried out by considering perturbative $\mathcal{R}^{2}$-corrections. However, there is no guarantee that this will still work at all orders in perturbation theory. In fact, the authors of Ooguri:2004zv argue —also providing some amount of evidence— that the exact black hole entropy should depend on the background hypermultiplet vevs. Crucially, though, the generalized prepotential (2.18) controlling the quantum entropy formula above is sensitive to the number of hypermultiplets but not to their vevs (see discussion after eq. (2.29) in the next section). Thus, from the macroscopic perspective it is not clear whether we could be missing some additional operators contributing to the black hole entropy, namely if (2.21) would be the end result of applying Wald’s procedure in the full Type IIA string theory. In Ooguri:2004zv a detailed analysis of the origin of (2.21) was performed and they suggested that it is computing instead a protected supersymmetric index. This idea has been supported by explicitly matching the black hole free energy [^7] This corresponds to the Legendre dual of the entropy and it gives the leading-order contribution to the gravitational path integral. with a supersymmetric index defined within the CFT living on the branes sourcing the BPS black hole background. In particular, the alternating signs of the terms which add up to give the index should account for the cancellation of the dependence on hypermultiplet vevs of the BPS states degeneracy. In what follows, we will not be concerned about whether (2.21) is truly computing an entropy or a protected supersymmetric index in Type IIA, and we will just focus on its properties along certain decompactification limits. With this subtlety in mind, we will refer to (2.21) simply as the BPS black hole entropy.
@@
It is also worth mentioning that in Ooguri:2004zv they revise the relation between macroscopic entropy and microstate counting performed in Strominger:1996sh . We present the subtlety following the modern review of Zaffaroni:2019dhb (see also references therein). What one should truly compute in order to compare the macroscopic and microscopic dual descriptions of the system is the partition function $\mathcal{Z}$. Such quantity can be formally defined via some path integral or as a microscopic generating function, respectively. In the latter case, it is not defined with a micro-canonical ensemble (i.e., with both electric $q$ and magnetic $p$ charges fixed), but rather with a mixed ensemble (fixed magnetic charges $p$ and electric potentials $\phi$). Thus, for a supersymmetric system, the partition function would have the form [^8] In the general case, there would also be some dependence on the angular momentum, which we omit here.
@@
with $F$ being some $\mathbb{Z}_{2}$-graded (i.e., fermionic) operator. The advantage of using the index is that it can also be evaluated from the macroscopic perspective. It is indeed an euclidean path integral with proper boundary conditions (which can be explicitly determined in concrete examples). In general, a supersymmetric index does not coincide with the partition function. However, in particular setups one can actually prove that the index provides a good approximation to the partition function. In essence, what one has to ensure is that there are no large cancellations between the different supersymmetric states over which we are tracing. [^10] In some examples this is automatically realized thanks to the symmetries exhibited by the configuration. In other case, the cancellation is avoided if the chemical potentials involved have complex phases. In general, though, there is not a unique, unambiguous prescription to find an appropriate $(-1)^{F}$ operator. What Ooguri:2004zv suggests then is that, despite (2.21) not being constructed as an index, it truly computes the Legendre dual of $\mathcal{Z}_{\text{index}}$ in the context of Type IIA string theory, and is therefore protected. Moreover, in the large charge expansion, we would also have $\mathcal{Z}_{\text{index}}\sim\mathcal{Z}$, so that it really captures the same quantity as $\mathcal{S}_{\text{micro}}$ and $\mathcal{S}_{\text{BH}}$, cf. eqs. (2.23) and (2.24).
@@ 2.3.1 Leading-order corrections to the generalized prepotential
$\mathcal{F}(X^{A})=$ $-\frac{1}{6}\mathcal{K}_{abc}\frac{X^{a}X^{b}X^{c}}{X^{0}}+K_{ab}^{(1)}X^{a}X^{b}+K_{a}^{(2)}X^{0}X^{a}+K^{(3)}(X^{0})^{2}$ (2.26)
@@
$\mathcal{K}_{abc}=\omega_{a}\cdot\omega_{b}\cdot\omega_{c}\,,\qquad K_{a}^{(2)}=\frac{1}{24}c_{2}(TX_{3})\cdot\omega_{a}\,,\qquad K^{(3)}=\frac{i\zeta(3)}{2(2\pi)^{3}}\chi_{E}(X_{3})\,$, (2.27)
whereas $K_{ab}^{(1)}$ can be fixed instead by requiring good symplectic transformation properties of the underlying period vector deWit:1992wf ; Harvey:1995fq . Similarly, the quantity $\chi_{E}(X_{3})=2\left(h^{1,1}(X_{3})-h^{2,1}(X_{3})\right)$ denotes the Euler characteristic of the threefold. Lastly, the coefficients $n_{\boldsymbol{k}}^{(0)}$ are known as genus-zero Gopakumar–Vafa invariants and count, for each positive homology representative $\boldsymbol{k}=k_{a}\gamma^{a}\in H_{2}^{+}(X_{3},\mathbb{Z})$, the indexed degeneracy of BPS D2-brane states wrapped on 2-cycles within the corresponding supersymmetric class Gopakumar:1998ii ; Gopakumar:1998jq .
On the other hand, the (holomorphic part of the) genus-1 topological string amplitude can be expanded around the large radius point as Bershadsky:1993ta ; Bershadsky:1993cx ; Katz:1999xq
$\mathcal{F}_{1}(X^{A})=\frac{1}{24}\int_{X_{3}}J_{\mathbb{C}}\wedge c_{2}(TX_{3})\,+\,\mathcal{O}\left(e^{2\pi iz^{a}}\right)=\frac{1}{24}c_{2,a}\,z^{a}\,+\,\mathcal{O}\left(e^{2\pi iz^{a}}\right)\,$, (2.28)
where $J_{\mathbb{C}}=z^{a}\,\omega_{a}=(b^{a}+it^{a})\,\omega_{a}$ is the complexified Kähler 2-form of the Calabi–Yau threefold, whilst $c_{2}(TX_{3})$ denotes the second Chern class of its tangent bundle. This contribution can be easily understood as coming from the dimensional reduction of an analogous four-derivative, curvature squared operator in 5d $\mathcal{N}=1$ supergravity Grimm:2017okk .
@@
where $n\in\mathbb{Z}$ is the 0-brane charge. Therefore, upon substituting this into (2.10) and performing the integral —taking account that each D0-brane yields $\frac{\chi(X_{3})}{2}$ times the contribution of a single hypermultiplet Gopakumar:1998jq — as well as the infinite sum, one finds (in units of $m_{\rm D0}/2\pi$)
@@ 2.3.2 Black hole solutions in the large volume patch
$q_{0}=\frac{i}{(Y^{0})^{2}}\left(D_{abc}Y^{a}Y^{b}Y^{c}+Y^{a}d_{a}\Upsilon\right)-i\,\frac{\partial G(Y^{0},\Upsilon)}{\partial Y^{0}}\ +\ \text{h.c.}\ .$ (2.37)
@@
and we isolate in (2.39) the terms depending only on $G(Y^{0},\Upsilon)$ together with its derivatives, we obtain the simpler formula [^11] Notice that, despite the piece $\mathcal{S}_{\rm BH}\big{\rvert}_{G=0}$ in (2.41) not depending explicitly on $G(Y^{0},\Upsilon)$, it still does implicitly via the stabilized value of $Y^{0}$, which also involves the higher-genus terms; cf. eq. (2.37).
@@ 3.1.1 The two-derivative analysis
with $C=e^{K/2}\bar{Z}$ some compensator field that ensures the symplectic and Kähler invariance of any solution to the attractor equations above (cf. discussion around (2.18)). Henceforth, we will concentrate on black holes characterized by having $p^{0}=0$, i.e., no D6-brane charge, as seen from the Type IIA perspective. Consequently, we deduce from (3.3) that the rescaled quantity $CX^{0}$ is purely real (and, e.g., non-negative), hence allowing us to completely solve the algebraic system as follows
@@
where $D^{ab}$ is the inverse matrix of $D_{ab}\equiv D_{abc}p^{c}$ —which is assumed to exist, [^12] This ensures that there exists a unique solution to the algebraic system (3.3), given precisely by (3.4). Notice that $D_{ab}$ could be singular in special circumstances, such that $D^{ab}$ might not be well-defined Marchesano:2023thx ; Marchesano:2024tod ; Castellano:2024gwi ; CMP . and we defined $\hat{q}_{0}\equiv q_{0}+\frac{1}{12}D^{ab}q_{a}q_{b}$. Note that this latter shift may be interpreted as an induced D0-brane charge in the worldvolume of the D2- and D4-branes comprising the 4d black hole of the form $\delta q_{0}=q_{a}\,\text{Re}\,z^{a}-3p^{a}D_{abc}\,\text{Re}\,z^{b}\,\text{Re}\,z^{c}$.
@@
$(\hat{q}^{0})^{2},\,\left(p^{a}\right)^{2}\gg\left|\frac{D_{abc}p^{a}p^{b}p^{c}}{\hat{q}_{0}}\right|\,$, (3.9)
@@ 3.2.1 Including higher-derivative corrections
$\mathcal{S}_{\rm BH}=-4\pi Y^{0}\hat{q}_{0}-i\pi\left(3Y^{0}G_{0}+2\Upsilon G_{\Upsilon}-\text{h.c.}\right)\,$, (3.14b)
@@
Finally, let us try to understand the extra constraints imposed by the hierarchy (3.12). The first one is required so as to ensure that the corrections to the attractor solution and black hole entropy due to the genus-1 term in (2.31) are in fact subleading with respect to the two-derivative results. The second condition, on the other hand, is more interesting, since its net effect is to suppress the higher-genus terms as well. [^13] Actually, it suppresses the full tower of quantum corrections $\delta\mathcal{F}_{g}$ associated to D0-brane states, as captured by $G(Y^{0},\Upsilon)$, which also includes a genus-0 contribution. Furthermore, it can be translated into an equivalent mathematical statement on the value for $y^{0}$ (equivalently $x^{0}$ in (3.4)), whose precise growth with the gauge charges has been ignored so far. To see this, we should first compute the imaginary part of $G_{0}(Y^{0},\Upsilon)$, which from the local 4d EFT perspective is given by the following asymptotic series
@@
$(\hat{q}^{0})^{2},\,\left(p^{a}\right)^{2}\gg\left|\frac{D_{abc}p^{a}p^{b}p^{c}}{\hat{q}_{0}}\right|\gg 1\,.$ (3.18)